# Improving Clinical Note Generation from Complex Doctor-Patient Conversation

### v1

@@ Abstract
Writing clinical notes and documenting medical exams is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually writing these notes is time-consuming and can impact the amount of time clinicians can spend on direct patient interaction and other tasks. Consequently, the development of automated clinical note generation systems has emerged as a clinically meaningful area of research within AI for health. In this paper, we present three key contributions to the field of clinical note generation using large language models (LLMs). First, we introduce CliniKnote, a comprehensive dataset consisting of 1,200 complex doctor-patient conversations paired with their full clinical notes. This dataset, created and curated by medical experts with the help of modern neural networks, provides a valuable resource for training and evaluating models in clinical note generation tasks. Second, we propose the K-SOAP (Keyword, Subjective, Objective, Assessment, and Plan) note format, which enhances traditional SOAP (Podder, Lew, and Ghassemzadeh 2024) (Subjective, Objective, Assessment, and Plan) notes by adding a keyword section at the top, allowing for quick identification of essential information. Third, we develop an automatic pipeline to generate K-SOAP notes from doctor-patient conversations and benchmark various modern LLMs using various metrics. Our results demonstrate significant improvements in efficiency and performance compared to standard LLM finetuning methods.
Introduction
The generation of clinical notes is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually writing these notes is time-consuming, often taking clinicians between 10 to 30 minutes per note. This extensive time requirement may impact the amount of time clinicians can spend on direct patient interaction and other tasks. Consequently, the development of automated clinical note generation systems has emerged as an important area of research within AI for health. Recent advancements in large language models (LLMs) have significantly enhanced text summarization capabilities, offering the potential to streamline the creation of clinical notes. These models can assist clinicians in efficiently drafting their notes, allowing them to focus more on patient care and manage multiple tasks during consultations (Knoll et al. 2022).
@@ Introduction
Despite this promise, the application in clinical note generation faces several challenges. Firstly, the quality of existing datasets is a major concern. While there are publicly available datasets like MTS-DIALOG  (Ben Abacha et al. 2023) and ACI-BENCH (wai Yim et al. 2023), they often fall short in terms of the complexity and length required to accurately simulate real-life clinical interactions. Secondly, the current format of clinical notes, such as SOAP notes, although clear and structured, can be time-consuming for clinicians to review, particularly as patient conditions become more complex. This increased complexity results in longer notes with dispersed critical information, complicating the decision-making process (Davidoff and Miglus 2011). Lastly, there is a notable lack of comprehensive benchmarking for LLM-based clinical note generation, hindering the evaluation and comparison of different models and approaches in this domain. Addressing these challenges requires finetuning LLMs with domain-specific knowledge, developing high-quality datasets that reflect real-life clinical dialogues, and establishing thorough benchmarking protocols to advance the efficacy and reliability of automated clinical note generation systems (Zhou et al. 2024).
@@
Our work makes several significant contributions to the field of clinical note generation using large language models. First, we introduce CliniKnote, a comprehensive dataset of complex doctor-patient conversations paired with full clinical notes. Second, we propose the K-SOAP note format, enhancing traditional SOAP notes with a keyword section to improve information retrieval. Third, we perform thorough benchmarking of various large language models on this task and develop a novel method that significantly improves over standard LLM finetuning approaches.
All conversations in CliniKnote are simulated and recorded by medical experts to ensure they closely resemble real-life doctor-patient interactions, and the paired notes are also written by real medical experts. This dataset provides a valuable resource for training and evaluating models in clinical note generation tasks.
@@
Using this dataset, we build the K-SOAP note format, which adds a keyword section at the top of traditional SOAP notes. This addition allows doctors to quickly recall domain knowledge at a glance. To achieve this, we combine clinical Named Entity Recognition (NER) and Relation Extraction (RE) to extract entities such as symptoms and diseases from the dialogue along with their relationships to the patient (e.g., present or absent). Previous clinical NER datasets like NCBI-disease (Dogan, Leaman, and Lu 2014) and BC5CDR (Li et al. 2016) only contain raw disease entities without relational information, highlighting the need for more comprehensive training resources like CliniKnote.
@@
Furthermore, we finetune various modern large language models and develop an automatic pipeline to generate K-SOAP notes from the conversations. The workflow of this pipeline is shown in figure 1. We benchmark CliniKnote using various automatic evaluation to assess the quality of the generated notes and their practical applicability in real-life clinical settings. We also investigate the effects of data augmentation, finetuned domain knowledge, and different adapters for various section segmentations. We found that domain knowledge finetuned models with 15 different adapters for each section achieve the best performance. Our results demonstrate that our proposed method significantly reduces time complexity and enhances the performance of automatically generating K-SOAP notes. The time required for reviewing and correcting notes generated by this pipeline is significantly shorter than the original reviewing and writing time.
Related Work
@@
Various methods have been developed for clinical note generation. For instance, Singh et al. (2023) build large-scale sequence-to-sequence models for generating clinical notes from patient-doctor conversations. The use of pre-trained large language models (LLMs) for this task is increasingly popular (Zhou et al. 2024). Biswas and Talukdar (2024) demonstrates advanced prompting techniques to generate draft clinical notes using LLMs, while Grambow, Zhang, and Schaaf (2022) introduces in-domain pre-training to enhance transformer models’ clinical summarization performance. Additionally, the MEDIQA-2023 Dialogue2Note shared tasks (Tang et al. 2023) include finetuning pre-trained dialogue summarization models and using few-shot in-context learning with GPT-4 to address medical dialogue summarization challenges.
@@
For publicly available clinical conversation-note datasets, there are several notable examples. Dr.Summarize (Joshi et al. 2020), generated from a telemedicine platform, and the dataset presented by Chintagunta et al. (2021), created using GPT-3, both follow a snippet-summary format. For well-structured SOAP medical notes, ACI-BENCH (wai Yim et al. 2023) and PriMock57 (Korfiatis et al. 2022) feature role-played dialogues with complete SOAP notes but contain only 207 and 57 data points, respectively, which is insufficient for training modern large language models. MTS-DIALOG (Ben Abacha et al. 2023), previously the largest dataset with clinical notes, includes 1,700 doctor-patient conversations (16k turns and 18k sentences) and summarized clinical notes (6k sentences). However, MTS-DIALOG consists of segmented snippets and paired notes that include only some sections rather than full notes.
@@
Named Entity Recognition (NER) and Relation Extraction (RE) are essential for extracting structured information from unstructured text in natural language processing. Recent advancements in deep learning, particularly with BiLSTM-CRF and transformer-based models like BERT, have significantly improved accuracy in identifying entities across various domains, including the medical field. Notably, recent work combining NER and RE for extracting complex information from scientific texts (Dunn et al. 2022) and utilizing label-supervised Llama finetuning (Li et al. 2023) has achieved state-of-the-art performance. Building on these advancements, our work integrates cutting-edge NER and RE techniques with large language models to enhance clinical note generation from doctor-patient dialogues.
As large language models scale up, finetuning and storing all the parameters is prohibitively costly and eventually becomes practically infeasible. Recent advancements in natural language processing have led to the development of parameter-efficient pre-trained large language models. These models aim to reduce the computational and memory requirements while maintaining or even improving performance across various tasks. One notable approach is the use of adapter modules, which allow for the finetuning of a small number of additional parameters while keeping the majority of the pre-trained model fixed (Houlsby et al. 2019). This technique has been proven effective in domain adaptation and multilingual settings. Additionally, methods such as LoRA (Low-Rank Adaptation) (Hu et al. 2021) and Qlora(Quantized Low-Rank Adaptation) (Dettmers et al. 2023) have shown promise in optimizing model efficiency by introducing minimal changes to the original model architecture. These innovations are particularly relevant for applications in the medical domain, where models need to handle large volumes of complex data without incurring prohibitive computational costs. Our work builds upon these foundations by integrating parameter-efficient techniques with state-of-the-art summarization and information extraction models to enhance the generation of clinical notes from doctor-patient dialogues.
@@
CliniKnote Dataset Assessment
In this section, we discusse how we synthesize the conversations, SOAP notes, and keyword sections to create our CliniKnote dataset. We also highlight CliniKnote’s advantages in terms of scale and structure by comparing with the previous largest public available note dataset MTS-DIALOGUE (Ben Abacha et al. 2023). The appendix A shows a full example of CliniKnote.
Data Synthesis
CliniKnote’s dialogue and SOAP note sections are curated in a similar process as described in  (Fareez et al. 2022). A team of composed of senior medical students, medical residents, and medical doctors conducted simulated medical conversations. Following the completion of each simulated medical exam, the individual assuming the role of the physician would write a standard SOAP clinical note summarizing the findings, assessment, and medical plan of the simulated visit. Because each exam was simulated, there is no personal health information included in the dataset. The audio of the conversation was initially transcribed using automatic speech recognition (ASR) and then manually validated and corrected to minimize word error rates (WER). Through manual validation the WER of the transcripts were estimated to be approximately 1%.
@@
Data Augmenting
During the development of medical exams, the simulations covered a limited number of allergies, thus the dataset lack diversity of the allergy field. Using vanilla CliniKnote for note generation might cause identical allergy generation ignoring the specific patient input. In order to enhance the data diversity of CliniKnote, we augment the allergy section of CliniKnote by randomly sampling more allergies. To be more specific, we first generate a list of 30 common allergies and pick a number from 0 to 4 randomly. Then we randomly pick that number of allergies from the list, feed those allergies and the original dialogue into GPT-4, and prompt GPT-4 to replace the original allergies with the allergies we feed. The prompt is shown in Appendix A After replacing the allergies, we do a string comparison to make sure that is the only modified part in the dialogue. Through data augmentation for the allergy part, we have a more complementary version of CliniKnote, which could be generalized for more clinical-related tasks.
@@
Data Statistics
The final CliniKnote we introduced has 1000 dialogues in training set and 200 dialogues in test set, both paired with the proposed K-SOAP medical notes including the sections of Chief Complaint, History Of Presenting Illness, Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication, Immunization History, Review of Systems, Vital Signs, Physical Exam, Diagnostic, Assessment, Plan, and keywords. In training set, We have 952,316 words and 489,594 words in medical notes. In testing set, we have 192,116 words in dialogues and 96,458 words in medical notes.
@@
Compared with MTS-DIALOG, we have more numerous data points in CliniKnote and more complex dialogues that are closer to real-life clinical conversations. Table 1 shows the comparison between our CliniKnote and MTS-DIALOG, which is the largest publicly available dataset of medical conversations and associated notes, demonstrating our large scales in word count and turns. Our word count is 5 times greater than that of MTS-DIALOG in dialogues and 7 times larger in summary. The final keyword section contains 12,286 short dialogues segmented from the original dialogues paired with the entities with relation labeled. Each short dialogue has 6 turns in total so that it will be easier for NER models to process. Table 2 shows the specific count for each type of entity.
K-SOAP Generation Based On Medical Dialogue
@@
The complete task aims to generate K-SOAP note based on the input dialogue. We divide this task into two parts. The first part is to generate the traditional SOAP note and the second is to do keyword extraction in NER format. The concatenation of these two parts made up to the final K-SOAP note.
SOAP Clinical Note Generation
@@
The main task for our CliniKnote dataset is to generate the traditional SOAP clinical notes. To compare the performance of models with different sizes, domain knowledge, and released versions, we selected the following publicly available models to finetune for our benchmarking: qCammel-13b (Toma 2023), a finetuned version of the Llama-2 13B model, trained on a distilled dataset of 15,000 medical domain instructions using QLoRA; Llama2-13b (Touvron et al. 2023); Llama3-8b (AI@Meta 2024); and OpenBioLLM-Llama3-8B (Ankit Pal 2024), which is finetuned on the biomedical domain. Additionally, we include commercial models such as GPT-3.5-turbo and GPT-4 for our benchmarking. We use offical QLoRA (Dettmers et al. 2023) to finetune all the large language models. For this supervised learning part, the input is the prompt filled with the original dialogue and the output is the clinical note in CliniKnote dataset reformatted in json mode as it will be more convenient to access. After finetuning, the Parameter-Efficient finetuning (PEFT) adapters (Mangrulkar et al. 2022) will be saved and loaded to do further evaluation.

### v2

@@ Abstract
Writing clinical notes is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually writing these notes is time-consuming and can impact the amount of time clinicians can spend on direct patient interaction. In this paper, we present three key contributions to the field of clinical note generation using large language models (LLMs). First, we introduce CliniKnote, a comprehensive dataset consisting of 1,200 complex doctor-patient conversations paired with their clinical notes. This dataset, created and curated by medical experts with the help of modern neural networks, provides a valuable resource for training and evaluating models in clinical note generation tasks. Second, we propose the K-SOAP (Keyword, Subjective, Objective, Assessment, and Plan) note format, which enriches traditional SOAP [23] (Subjective, Objective, Assessment, and Plan) notes by adding a keyword section at the top, allowing for quick identification of essential information. Meanwhile, these keyword information will be fed into LLMs as guidance for generating more accurate and comprehensive notes. Third, we develop an automatic pipeline to generate K-SOAP notes from doctor-patient conversations and benchmark modern LLMs using various metrics. Our results demonstrate significant improvements in efficiency and performance compared to standard LLM finetuning methods. Demo page: https://github.com/catalwaysright/K-SOAP-GOODLABS.git
Keywords:
Clinical Note Generation Large Language Model Named Entity Recognition Element-aware Summarization
1 Introduction
The generation of clinical notes is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually dealing with these notes is time-consuming, often taking clinicians 20 to 103 min per day authoring notes and 7 to 56 min per day viewing notes[13]. This extensive time requirement may impact the amount of time clinicians can spend on direct patient interaction and other tasks. Thus, our goal is to reduce the time for both writing and viewing the notes. Recent advancements in large language models (LLMs) have significantly enhanced text summarization capabilities, offering the potential to streamline the creation of clinical notes. These models can assist clinicians in efficiently drafting their notes, allowing them to focus more on patient care and manage multiple tasks during consultations [16].
@@ 1 Introduction
Despite this promise, the application in clinical note generation faces several challenges. Firstly, the quality of existing datasets is a major concern. While there are publicly available datasets like MTS-DIALOG  [3] and ACI-BENCH [33], they often fall short in terms of the complexity and length required to accurately simulate real-life clinical interactions. Meanwhile, we have LLM-generated dialogue datasets like Note-Chat [31], which do not have real medical experts involved in generating process. Secondly, the clinical note of complex doctor-patient conversations can be time-consuming for clinicians to review. This increased complexity results in longer notes with dispersed critical information, complicating the decision-making process [6]. Lastly, there is a notable need of comprehensive benchmarking for LLM-based clinical note generation, hindering the evaluation and comparison of different models and approaches in this domain. Addressing these challenges requires finetuning LLMs with domain-specific knowledge, developing high-quality datasets that reflect real-life clinical dialogues, and establishing thorough benchmarking protocols to advance the efficacy and reliability of automated clinical note generation systems [35].
@@
Our work makes several contributions to the field of clinical note generation using large language models. First, we introduce CliniKnote, a comprehensive dataset of complex doctor-patient conversations paired with full clinical notes and keywords. Second, we propose the K-SOAP note format, enriching traditional SOAP notes with a keyword section to improve information retrieval. Third, we perform thorough benchmarking of various large language models on two different but connected tasks (note generation and keyword detection) and develop a novel method that integrated notes with keywords and significantly improves over standard LLM finetuning approaches.
@@
All conversations in CliniKnote are simulated and recorded by medical experts to ensure they closely resemble real-life doctor-patient interactions, and the paired notes are also written by real medical experts. We build the K-SOAP note format, which adds a keyword section at the top of traditional SOAP notes. This addition allows doctors to quickly recall domain knowledge at a glance and provides great supplement when enhancing LLM note generation as an element-aware summarization [32]. To achieve this, we combine clinical Named Entity Recognition (NER) and Relation Extraction (RE) to extract entities such as symptoms and diseases from the dialogue along with their relationships to the patient (e.g., present or absent). Previous clinical NER datasets like NCBI-disease [8] and BC5CDR [18] only contain raw disease entities without relational information. There are like some similar works like clinical negation detection [20], which contains affirmative, negative, speculated and recommended clinical named entities, while in CliniKnote, more types of relations with named entities are introduced.
@@
Furthermore, we finetune various modern large language models and develop an automatic pipeline to generate K-SOAP notes from the conversations. The workflow of this pipeline is shown in figure 1. We benchmark CliniKnote using various automatic evaluation to assess the quality of the generated notes and their practical applicability in real-life clinical settings. We also investigate the effects of data augmentation, finetuned domain knowledge, and different adapters for various section segmentations. Our results show that the proposed method significantly reduces time complexity and improves the efficiency of generating K-SOAP notes, requiring much less time compared to manual note-writing.
2 Related Work
@@
Various methods have been developed for clinical note generation. For instance, Singh [27] builds large-scale sequence-to-sequence models for generating notes from patient-doctor conversations. Using LLM for clinical note generation is increasingly popular [35]. Biswas [4] demonstrates advanced prompting techniques to generate clinical notes using LLMs, while Grambow [12] introduces in-domain pre-training to enhance transformer models’ clinical summarization performance. Additionally, the MEDIQA-2023 Dialogue2Note shared tasks [28] include finetuning pre-trained dialogue summarization models and using few-shot in-context learning with GPT-4 to address medical dialogue summarization challenges.
@@
There are some notable public clinical dialogue-note datasets. Dr.Summarize [15], generated from a telemedicine platform, and the dataset presented by [5], created using GPT-3, both follow a snippet-summary format. For well-structured SOAP medical notes, ACI-BENCH [33] and PriMock57 [17] feature role-played dialogues with complete SOAP notes but contain only 207 and 57 data points, respectively, which is insufficient for training modern large language models. MTS-DIALOG [3], previously the largest dataset with clinical notes, includes 1,700 doctor-patient conversations (16k turns and 18k sentences) and summarized clinical notes (6k sentences). However, MTS-DIALOG consists of segmented snippets and paired notes that include only some sections rather than full notes.
@@
Named Entity Recognition (NER) and Relation Extraction (RE) are essential for extracting structured information from unstructured text in natural language processing. Many datasets in both fields are publicly available, recorded by BigBIO [11]. Recent advancements in deep learning, particularly with BiLSTM-CRF and transformer-based models like BERT, have significantly improved accuracy in identifying entities across various domains, including the medical field. Notably, recent work combining NER and RE for extracting complex information from scientific texts [9] and utilizing label-supervised Llama finetuning [19] has achieved state-of-the-art performance. Building on these advancements, our work integrates NER and RE with large language models to enhance clinical note generation from doctor-patient dialogues.
3 CliniKnote Dataset Assessment
In this section, we discusse how we synthesize the conversations, SOAP notes, and keyword sections to create our CliniKnote dataset. We also highlight CliniKnote’s advantages in terms of scale and structure by comparing with the previous largest public available note dataset MTS-DIALOGUE [3].
@@ 3 CliniKnote Dataset Assessment
3.1 Data Synthesis
CliniKnote’s dialogue and SOAP note sections are curated in a similar process as described in  [10]. A team of composed of senior medical students, medical residents, and medical doctors conducted simulated medical conversations. Following the completion of each simulated medical exam, the individual assuming the role of the physician would write a standard SOAP clinical note summarizing the findings, assessment, and medical plan of the simulated visit. Because each exam was simulated, there is no personal health information included in the dataset. The audio of the conversation was initially transcribed using automatic speech recognition (ASR) and then manually validated and corrected to minimize word error rates (WER). Through manual validation the WER of the transcripts were estimated to be approximately 1%.
@@ 3.1 Data Synthesis
3.2 Data Augmenting
During the development of medical exams, the simulations covered a limited number of allergies, and the occurrence of penicillin is 82.6% among the whole dataset. Thus the dataset lacks diversity of the allergy field. Using vanilla CliniKnote for note generation might cause identical allergy generation ignoring the specific patient input.
@@ 3.2 Data Augmenting
In order to enhance the data diversity of CliniKnote, we augment the allergy section of CliniKnote. To be more specific, we first generate a list of 30 common allergies suggested by Food Standard Agency(FSA) and pick a number from 0 to 4 randomly. Then we sample that number of allergies from the list, feed those allergies and the original dialogue into GPT-4, and prompt GPT-4 to replace the original allergies with the allergies we feed and maintain the consistency of context. After replacing the allergies, we do a string comparison to make sure that is the only modified part in the dialogue. Through data augmentation for the allergy part, we have a more complementary version of CliniKnote, which could be generalized for more clinical-related tasks.
3.3 Data Statistics
The final CliniKnote we introduced has 1000 dialogues in training set and 200 dialogues in test set, both paired with the proposed K-SOAP medical notes including the sections of Chief Complaint, History Of Presenting Illness, Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication, Immunization History, Review of Systems, Vital Signs, Physical Exam, Diagnostic, Assessment, Plan, and keywords.
@@ 3.3 Data Statistics
Compared with MTS-DIALOG, CliniKnote has more data points and more complex dialogues that are closer to real-life clinical conversations. Table 1 shows the comparison between CliniKnote and MTS-DIALOG, demonstrating our large scales in word count and turns. Word count is 5 times than that of MTS-DIALOG in dialogues and 7 times in summary. The final keyword section contains 12,286 short dialogues segmented from the original dialogues paired with the entities and relation labeled. Table 2 shows the specific count for each type of entity.
4 K-SOAP Generation Based On Medical Dialogue
@@
The complete task aims to generate K-SOAP note based on the input dialogue. We divide this task into two parts. The first part is to to do keyword extraction in NER format and the second is to generate the traditional SOAP note with the guide of keywords in two complex sections. The concatenation of these two parts made up to the final K-SOAP note.
4.1 Keyword Extraction
We generate the keyword section for each clinical dialogue to reduce the reviewing complexity and provide guidance for note generation. We treat this task as a NER task but with relation as the prefix. For keyword extraction, we use Label-supervised unmasked Llama finetuning [19]. Instead of autoregressive generation, we add a label predictor at the top of a Llama-7b model. The dialogue sequences decoded by Llama will be fed into this predictor to predict the labels. We also keep the idea of removing the causal masks because their empirical studies show that using token representations learned with causal masks significantly underperforms in token classification tasks. We merge the diseases and symptoms together into only 4 types, which are PRESENT, ABSENT, FAMILY and UNKNOWN. To be more specific, we set the word to tag dictionary as ‘O’: 0, ‘B-PRESENT’: 1, ‘B-ABSENT’: 2, ‘B-FAMILY’: 3, ‘B-UNKNOWN’: 4, ‘I-PRESENT’: 5, ‘I-ABSENT’:6, ‘I-FAMILY’:7, ‘I-UNKNOWN’:8 while B means start and I means inside. The model will determine whether a word is a valid entity and what is the relation with the patient at the same time.
4.2 SOAP Clinical Note Generation
@@
The main task for our CliniKnote dataset is to generate the traditional SOAP clinical notes. To compare the performance of models with different sizes, domain knowledge, and released versions, we selected the following public models to finetune for our benchmarking: qCammel-13b [29], a finetuned version of the Llama-2 13B model, trained on a distilled dataset of 15,000 medical domain instructions using Quantized Low-Rank Adaptation(QLoRA) [7]; Llama2-13b [30]; Llama3-8b [1]; and OpenBioLLM-Llama3-8B [2], which is finetuned on the biomedical domain. Additionally, we include commercial models such as GPT-3.5-turbo and GPT-4 for our benchmarking. We use offical QLoRA to finetune all the large language models. For the supervised learning, there are two versions of the input. The first consists of a prompt filled with the original dialogue, while the second enhances this prompt by incorporating keyword information. The output is the clinical note in CliniKnote dataset reformatted in json mode as it will be more convenient to access. After finetuning, the Parameter-Efficient finetuning (PEFT) adapters [22] will be saved and loaded to do further evaluation.
5 Evaluation of K-SOAP Note Generation
We evaluate CliniKnote using various metrics to assess note quality and time efficiency, highlighting its effectiveness in producing high-quality clinical notes and enhancing medical workflows.
5.1 Evaluation Metrics
For keyword evaluation, we only record and compare the accuracy, precision, recall and F1 score. For SOAP part note evaluation, we employ several metrics to ensure a comprehensive assessment. We start with the most traditional metric, ROUGE [21], which evaluates similarity based on common subsequences between the generated and reference texts. In addition to ROUGE, we introduce more advanced metrics: BERTScore [34] and BLEURT  [26]  [24]. These metrics leverage contextual embeddings from BERT to evaluate the semantic similarity between generated and reference texts. BERTScore uses token embeddings to compute precision, recall, and F1 scores, providing a more nuanced and context-aware evaluation compared to traditional metrics. Similarly, BLEURT incorporates pre-trained transformer models to better capture the meaning and fluency of the text, offering insights into the semantic quality of the generated notes. We use Roberta-large model for BERTScore metric and the newest checkpoint BLEURT-20 for BLEURT metric. We also include QuestEval [25] in our evaluation framework. Unlike traditional metrics, QuestEval does not require any ground truth references. Compared with ROUGE, QuestEval substantially improves the correlation with human judgments over four evaluation dimensions: consistency, coherence, fluency, and relevance. By incorporating question-answering models, QuestEval evaluates the generated text’s ability to provide accurate and relevant information, making it particularly suitable for complex domains like clinical note generation.
6 Experiments
Model Acc F1 Pre Recall Eval_Loss
biobert 0.933 0.735 0.755 0.715 0.197
scibert 0.921 0.669 0.723 0.622 0.221
pubmedbert 0.919 0.685 0.689 0.681 0.230
bluebert 0.903 0.554 0.679 0.467 0.285
bert 0.915 0.653 0.677 0.630 0.242
Llama2-7b-in-context - 0.692 0.391 0.500 -
Llama2-7b-QA - 0.642 0.731 0.657 0.307
Llama2-7b-
label-supervised 0.971 0.752 0.726 0.780 0.107
Table 3: Keyword performance comparison of different models.
6.1 Experimental Settings
We train all the models with one A100-80GB GPU. For clinical note generation task, we train all the large language models for two epochs. We use the default setting for question-answer finetuning with QLoRA. We set the learning rate to 0.0002 and the warmup ratio to 0.03. For QLoRA hyperparameters, we set LoRA $r$ to 64 and the loRA $\alpha$ to 16 and all the models are quantized in 4 bits. For keyword task, we finetune the model by LoRA. We set batch size as 8, learning rate as 1e-4, Max length as 256, LoRA $r$ as 256, LoRA $\alpha$ as 768, LoRA dropout as 0.1. We train Llama2-7b model for 5 epochs.
Keywords [‘1rash’, ‘|’, ‘1itchier’, ‘5at’, ‘5night’, ‘|’, ‘5sweat’]
[‘1rash’, ‘|’, ‘1itchier’, ‘5at’, ‘5night’]
Keywords [‘2cough’, ‘|’, ‘2shortness’, ‘6of’, ‘6breath’, ‘|’, ‘2changes’, ‘6in’, ‘6urination’, ‘|’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘|’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘|’, ‘2nausea’, ‘|’, ‘2vomiting’, ‘|’, ‘2abdominal’, ‘6pain’]
[‘2cough’, ‘|’, ‘2shortness’, ‘6of’, ‘6breath’, ‘|’, ‘2changes’, ‘6in’, ‘6urination’, ‘|’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘|’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘|’, ‘2nausea’, ‘|’, ‘2vomiting’, ‘|’, ‘2abdominal’, ‘6pain’]
6.2 Keyword Extraction Results
6.3 Clinical Note Generation Results
In this section, all models are finetuned and tested in 3 modes, which are full note mode, section-4 mode and section-15 mode that contain different numbers of LoRA adapter [14]. Full note mode means we only train one adapter to generation the full clinical note. Section-4 mode means we train four different adapters to generate the full note. How we divide the four parts is a little different from how we separate Subjective, Objective, Assessment, and Plan parts. We only have Chief Complaint and History Of Presenting Illness in the first part. The second part contains Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication List, Immunization History and Review of Systems. The third part contains Vital Signs and Physical Exam while the final part contains Assessment, Diagnosis and Plans. Section-15 mode means we train 15 different adapters to generate all the fifteen sections contained in the note separately. For each section, we evaluate the performance difference of our best model with and without the inclusion of keyword guidance. As for commercial models, we try full note zeroshot generation and oneshot(one demo in prompt) in-context learning. The total generation time is all below 2 min.